OpenStudy (anonymous):

Question 1: What is the probability of rolling an even number on a die and then flipping a coin and having it land on heads? Question 2: A bag contains 8 white marbles, 5 blue marbles, and 3 red marbles. What is the probability of choosing a white marble, not replacing it, and then choosing a red marble?

OpenStudy (anonymous):

Question 1 = 8.3% I believe

OpenStudy (kirbykirby):

There are three ways to get an even number: 2, 4, 6 Thus, the probability of getting an even number is \(3/6=1/2\) Probability of getting heads = 1/2 So, the probability P(even number, heads) = P(even number) * P(heads) = 1/2 * 1/2 = 1/4

OpenStudy (kirbykirby):

2nd question: 1st draw: probability of choosing white = 8/16 = 1/2 (since there are 8 white marbles, and 16 in total) 2nd draw: you took a white one already, so there are 7 white marbles left, and thus 15 in total, so the probability of getting red is thus 3/15 = 1/5 so the total probability is: 1/2 * 1/5 = 1/10